Write an equation of the line that passes thru the given points:
(3,-12), (8,4)
You have to find the slope intercept first right?
Then use y-y1=m(x-x1) right?
Or am I way off
Thanks

Answer by checkley71(8403)

(y2-y1)/(x2-x1) = slope
(4+12)/(8-3)
16/5 is the slope now substitute one set of points in the line equation thus
Y=mX+b
-12=16/5*3+b
-12=48/5+b
b=-12-48/5
b=(-60-48)/5
b=-108/5 OR -21 3/5 SO THE LINE EQUATION IS
Y=16/5X-108/5